OpenStudy (anonymous):

Need help with log..1 question..giving out medals

OpenStudy (anonymous):

you just want to know what log (1) = ?

OpenStudy (anonymous):

please help

zepdrix (zepdrix):

\[\Large 2\log_5x+\frac{1}{2}\log_5y-4\log_5z=\log_5\frac{x^2\sqrt{y}}{z^4}\] So we'll need to use a few rules of logarithms. \[\large \color{royalblue}{b\cdot\log(a)=\log(a^b)}\]We'll use this rule to deal with the coefficients in front of the logs.

OpenStudy (kainui):

Just for the record, log(1) is not a part of this problem.

zepdrix (zepdrix):

\[\Large 2\log_5x=\log_5x^2\]Understand how that works using our rule?

OpenStudy (anonymous):

but how do I make it the same?

zepdrix (zepdrix):

We're doing that step by step.

OpenStudy (anonymous):

I was referring to 1 question not log 1

OpenStudy (anonymous):

oh ok..can you show me?

zepdrix (zepdrix):

I just showed you the first step above, do you understand so far?

OpenStudy (anonymous):

yes I do

zepdrix (zepdrix):

\[\Large \frac{1}{2}\log_5y\qquad=\qquad \log_5y^{1/2}\]We can rewrite this using root notation,\[\Large \log_5\sqrt y\] So using this first rule, what we have so far is,\[\Large \log_5x^2+\log_5\sqrt y-4\log_5z=\log_5\frac{x^2\sqrt{y}}{z^4}\] Can you follow this idea to simplify this third term?\[\Large 4\log_5z \qquad = \qquad ?\]

OpenStudy (anonymous):

log 5z^4?

zepdrix (zepdrix):

\[\Large \log_5x^2+\log_5\sqrt y-\log_5z^4=\log_5\frac{x^2\sqrt{y}}{z^4}\]Ok good! From here we'll need to apply two more rules of logs:\[\large \color{green}{\log(a)+\log(b)=\log(ab)}\]\[\large \color{purple}{\log(b)-\log(c)=\log\left(\frac{b}{c}\right)}\] When we have the sum of logs, we can rewrite them as a single log with the product of their arguments inside. Applying this to our first two terms gives us,\[\Large \log_5x^2+\log_5\sqrt y\qquad=\qquad \log_5\left(x^2\sqrt y\right)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is next?

OpenStudy (kainui):

What do you think comes next?

zepdrix (zepdrix):

\[\Large \log_5\left(x^2\sqrt y\right)-\log_5z^4=\log_5\frac{x^2\sqrt{y}}{z^4}\] So next we'll want to combine these two terms,\[\Large \log_5\left(x^2\sqrt y\right)-\log_5z^4 \qquad=\qquad ?\] using the purple rule I listed.

OpenStudy (kainui):

You should really think about this, because at this point you only have one step and it is very similar to the step he just showed you.

OpenStudy (anonymous):

log 5 (x^2 qrt y)/ log 5z^4

OpenStudy (anonymous):

thank you!

zepdrix (zepdrix):

Close! We don't want to divide logs,\[\Large \log_5\left(x^2\sqrt y\right)-\log_5z^4 \qquad\ne\qquad \frac{\log_5\left(x^2\sqrt y\right)}{\log_5z^4}\] we want to divide the `insides`. \[\Large \log_5\left(x^2\sqrt y\right)-\log_5z^4 \qquad=\qquad \log_5\left(\frac{x^2\sqrt y}{z^4}\right)\]

OpenStudy (anonymous):

thank you